If the perimeter of a rectangle is 50 cm and the area is 156 square cm, then what are the dimensions of the rectangle?

The width, or shorter side is___ cm

The length, or longer side is ___ cm
Answer #1

Solution : L-length b b- breadth perimeter of rectangle = 22+2b = soom Area of Rectangle 16 = 156 cm? 22 50-2b i = 50-26 2 su